The objective of the controller design in Fig. 1 is to find the controller Gc(s) such that I.)The closed-loop system is stable ii.)The output of the system above (y(t)) can track the reference input r(t) = At (A>0 is any real number). Use the Nyquist plot and Nyquist criterion to show that: a.)The portional controller Gc(s)=K can achieve asymptotic tracking of the ramp input r(t)=At but cannot meet the stability requirement for 0<K<+inf. b.) Can the PI controller Gc(s)=K(1+10/s) be used to meet the requirements I.) - ii.)? Please use the Nyquist plot and Nyquist criterion, together with the steady-state analysis, to prove your conclusion. c.) Can the PI controller Gc(s)=K(1+10/s) be used to track the input signal r(t)=At^2 for any A>0? Prove or disprove your answer.plant [ problem 1 CPI compensation ) RS) +9+40 stopy Controller Fig. I The objective of the controller design in Fig. 1 is to
Answer #1

1)

S+10 S2 Controller plant let GeCS)= kp wat kpisto s? fps + 10kp KPS+ 82 = $? + Kopisto) RIS) Itke (st10 1 stable for all clos

2)

Alyqurile yu) kisho) we draw Nyquist plot for Oltr ie open loop Transfer function of al system put Sujuvuelo 1 (100+10) - Mag

3)

O QC3) = K (H+.10/3) = *( + 10) x (stre) - K(SH10)2 yes) K (840)2 53 - K(s+10)2 ROD SP+ x(st 2012 5P+k[s+206 +100] = K (stid?

4)

© g-tolt): At? | Ges) = ( 14 iols) Res) - 42 $ 3 wikit 11) - Gels). (6+10) K(Sto/S+10) ROS) it Gers) (Strol HKISHO)? S3 yeo)

SUMMARY: here we have to find out characteristic equation of closed loop trnsfer function and cheack for stability before continuing furthur analysis we must ensure weather the system is stable or not.We have to apply laplase trasform to convert time domain input into s domain.

trick:while drawing nyquist plot you can follow these three simple steps

i)draw polar plot

ii)draw its mirror image

iii)join the mirror image end direction with polar plot starting direction

HOPE THIS HELPS YOU!!

HAPPY CHEGGING!!

*you can encourage my efforts by giving a upvote.